\begin{thebibliography}{Wah81}

\bibitem[Bri68]{Brieskorn-1967-1968}
E. Brieskorn.
\newblock Rationale {S}ingularit\"aten komplexer {F}l\"achen.
\newblock {\em Invent. Math.}, \textbf{4}:336--358, 1967/1968.

\bibitem[Dur79]{Durfee1979}
A. Durfee.
\newblock Fifteen characterizations of rational double points and simple
  critical points.
\newblock {\em Enseign. Math. (2)}, 25:132--163, 1979.

\bibitem[Hac04]{Hacking2004}
P. Hacking.
\newblock Compact moduli of plane curves.
\newblock {\em Duke Math. J.}, \textbf{124}(2):213--257, 2004.

\bibitem[Hac12]{Hacking2012}
P. Hacking.
\newblock Compact moduli spaces of surfaces of general type.
\newblock In {\em Compact moduli spaces and vector bundles. Conference on
  compact moduli and vector bundles, University of Georgia, Athens, GA, USA,
  October 21--24, 2010}, pages 1--18. Providence, RI: American Mathematical
  Society (AMS), 2012.

\bibitem[Hac13]{Hacking2013}
P. Hacking.
\newblock Exceptional bundles associated to degenerations of surfaces.
\newblock {\em Duke Math. J.}, 162(6):1171--1202, 2013.

\bibitem[HP10]{Hacking-Prokhorov-2010}
P. Hacking and Y. Prokhorov.
\newblock Smoothable del {P}ezzo surfaces with quotient singularities.
\newblock {\em Compositio Math.}, \textbf{146}(1):169--192, 2010.

% \bibitem[HTU13]{Hacking-Tevelev-Urzua2013}
% P. Hacking, J. Tevelev, and G. Urz\'ua.
% \newblock Flipping surfaces.
% \newblock {\em ArXiv e-print}, 1310.1580, 2013.


\bibitem[KN98]{Karpov1998}
B. Karpov and D. Nogin.
\newblock Three-block exceptional sets on del {P}ezzo surfaces.
\newblock {\em Izv. Ross. Akad. Nauk Ser. Mat.}, \textbf{62}(3):3--38, 1998.

\bibitem[KM98]{Kollar-Mori-19988}
J. Koll{\'a}r and S. Mori.
\newblock {\em Birational geometry of algebraic varieties.} {\em
  Cambridge Tracts in Mathematics} \textbf{134}.
\newblock Cambridge University Press, Cambridge, 1998.

\bibitem[KSB88]{Kollar-ShB-1988}
J.~Koll{\'a}r and N. Shepherd-Barron.
\newblock Threefolds and deformations of surface singularities.
\newblock {\em Invent. Math.}, \textbf{91}(2):299--338, 1988.


\bibitem[Kol92]{Utah}
J.~Koll{\'a}r, editor.
\newblock {\em Flips and abundance for algebraic threefolds}.
\newblock Soci\'et\'e Math\'ematique de France, Paris, 1992.
\newblock Papers from the Second Summer Seminar on Algebraic Geometry held at
  the University of Utah, Salt Lake City, Utah, August 1991, Ast\'erisque 
  \textbf{211} (1992).



\bibitem[Man91]{Manetti-1991}
M. Manetti.
\newblock Normal degenerations of the complex projective plane.
\newblock {\em J. Reine Angew. Math.}, \textbf{419}:89--118, 1991.

\bibitem[Mor85]{Morrison-1985}
D. Morrison.
\newblock The birational geometry of surfaces with rational double points.
\newblock {\em Math. Ann.}, \textbf{271}(3):415--438, 1985.

\bibitem[MZ88]{Miyanishi-Zhang-1988}
M.~Miyanishi and D.-Q. Zhang.
\newblock Gorenstein log del {P}ezzo surfaces of rank one.
\newblock {\em J. Algebra}, \textbf{118}(1):63--84, 1988.

\bibitem[Pro01]{Prokhorov-2001}
Y. Prokhorov.
\newblock {\em Lectures on complements on log surfaces}. {\em MSJ
  Memoirs} \textbf{10}.
\newblock Mathematical Society of Japan, Tokyo, 2001.

\bibitem[Pro04]{Prokhorov-2004b}
Y. Prokhorov.
\newblock On semistable {M}ori contractions.
\newblock {\em Izv. Ross. Akad. Nauk Ser. Mat.}, \textbf{68}(2):147--158, 2004.

\bibitem[PS09]{Prokhorov-Shokurov-2009}
Y. Prokhorov and V. Shokurov.
\newblock Towards the second main theorem on complements.
\newblock {\em J. Algebraic Geom.}, \textbf{18}(1):151--199, 2009.

\bibitem[Rei87]{Reid-YPG1987}
M. Reid.
\newblock Young person's guide to canonical singularities.
\newblock In {\em Algebraic geometry, Bowdoin, 1985 (Brunswick, Maine, 1985)},
{\em Proc. Sympos. Pure Math.} \textbf{46},  345--414. Amer. Math.
  Soc., Providence, RI, 1987.

\bibitem[Sho00]{Shokurov-2000}
V. Shokurov.
\newblock Complements on surfaces.
\newblock {\em J. Math. Sci. (New York)}, \textbf{102}(2):3876--3932, 2000.


\bibitem[Wah81]{Wahl-1981}
J. Wahl.
\newblock Smoothings of normal surface singularities.
\newblock {\em Topology}, \textbf{20}(3):219--246, 1981.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{mathrsfs, amssymb}
\usepackage[matrix,arrow,curve]{xy}

\makeatletter \@addtoreset{equation}{section} \makeatother

\newcommand{\FF}{\mathbb{F}}

\newcommand{\CC}{\mathbb{C}}
\newcommand{\QQ}{\mathbb{Q}}

\newcommand{\ZZ}{\mathbb{Z}}
\newcommand{\PP}{\mathbb{P}}

\newcommand{\OOO}{{\mathscr{O}}} 
\newcommand{\comp}{{{\,\scriptstyle{\circ}\,}}}
\newcommand{\down}[1]{\left\lfloor #1\right\rfloor}
\newcommand{\fr}[1]{\left\{ #1\right\}}

\newcommand{\muu}{{\boldsymbol{\mu}}}

\newcommand{\lin}{\text{---}}

\newcommand{\A}{\operatorname{A}}
\newcommand{\D}{\operatorname{D}}
\newcommand{\E}{\operatorname{E}}

\newcommand{\xref}[1]{{\rm \ref{#1}}}

\newcommand{\s}{\operatorname{s}}
\newcommand{\Pic}{\operatorname{Pic}}
\newcommand{\Sing}{\operatorname{Sing}}

\newcommand{\Bs}{\operatorname{Bs}}
\newcommand{\red}{\operatorname{red}}
\newcommand{\Supp}{\operatorname{Supp}}
\newcommand{\NE}{\overline{\operatorname{NE}}}
\newcommand{\Diff}{\operatorname{Diff}}
\newcommand{\T}{$\operatorname{T}$}

\renewcommand{\emptyset}{\varnothing}

\newtheorem{mtheorem}[equation]{}
\newtheorem{stheorem}{}[equation]
\newtheorem*{claim*}{Claim}

\theoremstyle{definition}
\newtheorem{mdefinition}[equation]{}
\newtheorem{sdefinition}[stheorem]{}

\title{A note on degenerations of del Pezzo surfaces}

\author{Yuri Prokhorov}
\thanks{
Partially supported by RScF grant no. 14-21-00053.
}

\address{
Steklov Institute of Mathematics, 8 Gub2kina street, Moscow 119991, Russia
\newline\indent
Department 
of Algebra, Faculty of Mathematics, Moscow State
University, Moscow 117234, Russia
\newline\indent
Laboratory of Algebraic Geometry, SU-HSE, 
7 Vavilova Str., Moscow 117312, Russia
}
\email{prokhoro@gmail.com}

\begin{document}

\maketitle
\section{Introduction}
\begin{mdefinition} 
This paper continues the classification of $\QQ$-Gorenstein degenerations of del Pezzo surfaces 
started in \cite{Manetti-1991}, \cite{Hacking-Prokhorov-2010}.
Let $\mathcal X\to \mathcal Z$ be a family of surfaces over a smooth curve $\mathcal Z$
such that a general fiber is a smooth del Pezzo surface and the special fiber 
$X:=\mathcal X_o$ is reduced, normal and has only quotient singularities.
Assume further that $\mathcal X$ is $\QQ$-Gorenstein and $-K_{\mathcal X}$
is ample over $\mathcal Z$. Such kind of families 
appear naturally in the three-dimensional minimal model program \cite{Kollar-ShB-1988}, \cite{Kollar-Mori-19988}
and in the study of certain moduli spaces  \cite{Hacking2004}, \cite{Hacking2012}.
It expected that the combinatorial structure of singularities of $X$
is related to exceptional vector bundles on smooth del Pezzo surfaces 
but this relation still looks mysterious (cf. \cite{Hacking2013}).

In this paper study  the special fiber $X$ of the above family 
under the condition that the Picard number $\rho(X)$ is large.
The case $\rho(X)=1$ was investigated completely in \cite{Hacking-Prokhorov-2010}.
Our main result is the following.
\end{mdefinition}

\begin{mtheorem}{\bf Theorem.} \label{main}
Let $X$ be a del Pezzo surface with only quotient singularities 
and let $\s(X)$ be the number of its non-Du Val points\footnote{For various definitions of Du Val (rational double) 
singularities we refer to 
\cite{Durfee1979}.}.
Assume that $X$ admits a $\QQ$-Gorenstein smoothing. Then 
\begin{enumerate}
\item 
$\s(X)\le \rho(X)+2$,
\item 
if $\s(X)=\rho(X)+2$, then $X$ is toric,
\item 
if $\s(X)=\rho(X)+1$, then $X$ admits an effective $\CC^*$-action.
\end{enumerate}
\end{mtheorem}

Similar to  \cite[Th. 1.3]{Hacking-Prokhorov-2010},
as a consequence of our techniques we verify a particular case 
of Reid's general elephant conjecture (cf.\cite[3.4B]{Reid-YPG1987}):

\begin{mtheorem}{\bf Theorem.}\label{thm-ge}
Let $f\colon \mathcal{X}\to \mathcal{Z}$ be a del Pezzo fibration over a smooth curve.
That is, $\mathcal{X}$ is a $3$-fold with terminal singularities, 
$f$ has connected fibers, and $-K_\mathcal{X}$ is ample over $\mathcal{Z}$.
Fix a point $o\in \mathcal{Z}$ and assume that the fiber $f^{-1}(o)$ is reduced, irreducible, normal, 
and has only quotient singularities.
Then, for some ample divisor $A$ on $\mathcal{Z}$, a general member $S\in |-K_\mathcal{X}+f^*A|$ 
is normal and has only Du Val singularities in a neighborhood of $f^{-1}(o)$.
\end{mtheorem}

Furthermore, we give a characterization of log surfaces that admit a $\CC^*$-actions
(Theorem \ref{theorem-toric})
and   establish the existence of $1$-complements 
on arbitrary del Pezzo surfaces with \T-singularities (Theorem \ref{th-1-compl}).
\par\medskip\noindent
\textbf{Acknowledgments.}

The work was written during my stay at the
International Centre for Theoretical Physics, Trieste.
I would like to thank ICTP for hospitality and support.

\section{Preliminaries}
\label{section-Preliminaries}
\begin{mdefinition}
Throughout this paper, we work over the field $\CC$ of complex numbers.
$\rho(X)$ denotes the Picard group of a variety $X$.
We use standard definitions, notation, and facts of the Minimal Model Program
\cite{Utah}, \cite{Kollar-Mori-19988}.

\begin{sdefinition}{\bf Proposition-Definition \cite[3.6]{Reid-YPG1987}, \cite[5.19]{Kollar-Mori-19988}.}
\label{def-canonical-index-one-covering}
Let $P\in X$ be the germ of a normal variety at a point $P$ and let $D$ be 
a Weil divisor on $X$. 
Assume  that $D$ is $\QQ$-Cartier at $P$, that is, $rD$ is
a Cartier divisor near $P$ for some positive integer $r$. 
Suppose that $r$ is the smallest such $r$. Then $r$ is called the \textit{index} of $D$.

There exists a covering $\pi: X^\sharp\to X$ which is Galois with 
group $\muu_r$ such that $ X^\sharp$ is normal, $\pi$ is etale over 
the the locus $X_0\subset X$ where $K_X$ is Cartier, and $P^\sharp:=\pi^{-1}(P)$ is a single point.
The divisor $K_{X^\sharp}=\pi^*K_X$ is Cartier.
Such a covering is called \textit{canonical index-one covering} of $P\in X$.
\end{sdefinition}
\end{mdefinition}

\begin{mdefinition}{\bf \T-singularities \cite{Kollar-ShB-1988}, \cite{Wahl-1981}.}
\label{Tdefclass}

\begin{sdefinition}{\bf Definition.}
Let $X$ be a normal surface such that $K_X$ is $\QQ$-Cartier.
We say that a deformation $\mathcal X/(0 \in S)$ over a germ $(0 \in S)$ with $0$-fiber $\mathcal X_0=X$ 
is \emph{$\QQ$-Gorenstein} if, locally analytically at every point $P \in X$,
$\mathcal X/S$ is induced by an equivariant deformation of 
the canonical index-one covering $(X^\sharp\ni P^\sharp)\to (X\ni P)$
(see \ref{def-canonical-index-one-covering}).
\end{sdefinition}

\begin{sdefinition}{\bf Definition \cite[Def.~3.7]{Kollar-ShB-1988}.} \label{defT}
Let $P \in X$ be a quotient singularity of dimension $2$.
We say $P \in X $ is a \emph{\T-singularity} if it admits a $\QQ$-Gorenstein smoothing.
That is, there exists a $\QQ$-Gorenstein deformation of $P \in X$ over a smooth curve germ
such that the general fiber is smooth.
\end{sdefinition}

\begin{stheorem}{\bf Proposition \cite[Prop.~3.10]{Kollar-ShB-1988}.} 
\label{Tclass}
A \T-singularity is either a Du~Val singularity or a cyclic quotient singularity of the form 
$\frac{1}{dn^2}(1,dna-1)$ for some positive integers $d$, $n$, $a$ with $\gcd(a,n)=1$.
\end{stheorem}
\end{mdefinition}

\begin{mdefinition}{\bf Noether's formula.}
For a \T-singularity $P \in X$, define
\[
\mu_P= 
\begin{cases}
r& \text{if $P \in X$ is a Du Val singularity of type $\A_r$, $\D_r$, or $\E_r$},\\
d-1&\text{if $P \in X$ is of type $\frac{1}{dn^2}(1,dna-1)$}. 
\end{cases}
\]
This number coincides with 
the \emph{Milnor number} of $P \in X$
\cite[Sec.~3]{Manetti-1991}.

\begin{stheorem}{\bf Proposition \cite{Hacking-Prokhorov-2010}.} \label{noether}
Let $X$ be a projective rational surface with \T-singularities. Then 
\[
K_X^2+\rho(X) + \sum_{P \in \Sing X} \mu_P = 10.
\]
\end{stheorem}

\begin{stheorem}{\bf Corollary.}\label{defspreservingrho}
Let $X$ be a del Pezzo surface with \T-singularities.
Then $K_X^2+\rho(X)\le 10$.
\end{stheorem}
\end{mdefinition}

\begin{mdefinition}{\bf Del Pezzo surfaces with \T-singularities.}
\begin{stheorem}{\bf Proposition \cite{Hacking-Prokhorov-2010}.} \label{dimension}
Let $X$ be a projective surface with \T-singularities such that $-K_X$ is nef and big. Then 
\[
\dim H^0(X,\OOO_X(-nK_X))= \frac12 n(n+1) K_X^2+1.
\]
\end{stheorem}
 \begin{stheorem}{\bf Corollary.} \label{dimension1}
Let $X$ be a projective surface with \T-singularities such that $-K_X$ is nef and big. Then 
\[
\dim |-K_X|= K_X^2>0.
\]
\end{stheorem}
\end{mdefinition}

\begin{mdefinition}{\bf Divisorial adjunction \cite[ch. 16]{Utah}.}
\label{def-diff}
Let $X$ be a normal variety and $S\subset X$ a reduced subscheme
of pure codimension one. Assume that the pair $(X,S)$ is
lc (log canonical \cite[2.10]{Utah}) in codimension two. Then there exists a naturally defined effective
$\QQ$-Weil divisor $\Diff_S(0)$, called the \emph{different},
such that
\[
(K_X+S)|_S=K_S+\Diff_S(0).
\]
Now let $B$ be a $\QQ$-divisor, which is $\QQ$-Cartier in
codimension two. Then the different for $K_X+S+B$ is defined by
the formula
\[
(K_X+S+B)|_S=K_S+\Diff_S(B). 
\]
In particular, if $B$ is a boundary and $(X,S+B)$ is lc in
codimension two, then $B$ is $\QQ$-Cartier in codimension two. Moreover, none of the
components of $\Diff_S(B)$ are contained in the singular locus of
$S$.
\end{mdefinition}

\begin{mdefinition}{\bf Classification of two-dimensional log canonical pairs
with reduced boundary \cite[ch. 3 \& Prop. 16.6]{Utah}, \cite[Th. 4.15]{Kollar-Mori-19988}.}
\label{2-dim-lc-pairs}
Let $P\in (X,C)$ be the germ of a two-dimensional log pair where 
$X$ is normal and $C$ is a (possibly reducible) reduced curve.
Assume that $(X,C)$ is lc. Then  one of the
following possibilities holds where all isomorphisms are 
isomorphisms of analytic germs:

\begin{sdefinition}\label{2-dim-lc-pairs-plt}
{\bf $(X,C)$  is plt {\rm (purely log terminal, \cite[2.13]{Utah}).}} Then 
\[
(X, C)\simeq(\CC^2, \{x_1=0 \})/\muu_{m}(1, a),\quad\text{with}\quad
\gcd(a, m)=1,\ m\ge 1,
\]
\[
\operatorname{Index}(K_X+C) =\operatorname{Index}(C) = m, \quad
\Diff_C(0)=(1-1/m)P.
\]
\end{sdefinition}

\begin{sdefinition}\label{2-dim-lc-pairs-lc-D}
{\bf $(X,C)$  is not plt and $C$ analytically reducible.} Then
\[
(X, C)\simeq(\CC^2, \{x_1x_2=0 \})/\muu_{m}(1, a),
\quad\text{with}\quad\gcd(a, m)=1,\ m\ge 1,
\]
\[
\operatorname{Index}(K_X+C) =1, \quad
\Diff_C(0)=0.
\]
\end{sdefinition}

\begin{sdefinition}\label{2-dim-lc-pairs-lc}
{\bf $(X,C)$  is not plt and $C$ analytically irreducible.} Then
\[
(X, C)\simeq(\CC^2, \{x_1x_2=0 \})/G,
\]
\[
\operatorname{Index}(K_X+C) =2, \quad
\Diff_C(0)=1,
\]
where $G\subset \operatorname{GL}_2(\CC)$ is a finite subgroup of dihedral type
without reflections (see \cite{Brieskorn-1967-1968} for the precise description
of $G$). 
\end{sdefinition}
\end{mdefinition}

\begin{mdefinition} {\bf $1$-complements.}
Let $X$ be a normal variety and let $D$ be a boundary on $X$
(an effective $\QQ$-divisor with coefficients $\le 1$).
Write $D=S+B$, where $S:=\down D$ (resp. $B:=\fr{D}$) is the integral 
(resp. fractional) part of $D$. A \emph{$1$-complement}
of $K_X+D$ is a divisor $D^+ \in |-K_X|$ such that $(X,D^+)$ is log canonical and 
$D^+ \ge S+\down{2B}$. In particular, if $D=0$, then a $1$-complement of 
$K_X$ is a divisor $D^+ \in |-K_X|$ such that $(X,D^+)$ is log canonical.
We say that 
the log divisor $K_X+D$ is \emph{$1$-complementary} if there exists a $1$-complement of $K_X+D$.

\begin{stheorem}{\bf Proposition  \cite[Prop. 4.3.2]{Prokhorov-2001}.}
\label{bir-prop}
Let $f\colon X\to Y$ be a birational contraction
and let $D$ be a boundary on $X$ such that
\begin{enumerate}
 \item 
\label{prop-i}
$K_X+D$ is nef over $Y$;
 \item 
\label{prop-ii}
the coefficients of $f_*D$ satisfy the
inequality $d_i\ge 1/2$.
\end{enumerate}
Assume that $K_Y+f_*D$ is $1$-complementary. Then so is $K_X+D$.
\end{stheorem}

\begin{stheorem}{\bf Proposition  \cite[Prop. 4.4.1]{Prokhorov-2001}.}
\label{prodolj}
Let $(X,D=S+B)$ be a log variety, where
$S:=\down{D}$ and $B:=\fr{D}$. Assume that
\begin{enumerate}
 \item 
$K_X+D$ is plt;
 \item 
$-(K_X+D)$ is nef and big;
 \item 
$S\ne 0$;
 \item 
\label{prop-2-iv}
the coefficients of $D=\sum d_iD_i$ satisfy the inequality $d_i\ge 1/2$.
\end{enumerate}
Further, assume that  there exists a
$1$-complement $K_S+\Diff_ S(B)^+$ of $K_S+\Diff_ S(B)$. Then 
 there exists a $1$-complement $K_X+S+B^+$ of
$K_X+S+B$ such that $\Diff_S(B)^+=\Diff_ S(B^+)$.
\end{stheorem}
\end{mdefinition}

\begin{mdefinition}{\bf Contractions of surfaces with  Du Val singularities.}
For convenience of the reader we state facts about MMP for Du Val surfaces.

\begin{sdefinition}{\bf Definition (see \cite{Morrison-1985}).}\footnote{For general definition see e.g. 
\cite[4.56]{Kollar-Mori-19988} or \cite[3.2]{Prokhorov-2001}.}
Let $y\in Y$ be a smooth point on a surface and let $(u,v)$ be a 
local coordinates near $y$.
A \textit{weighted blowup} with weights $(1,n)$ of a $y\in Y$
is the blowup $X\to Y$ of the ideal $(u,v^{n})$.
\end{sdefinition}

Clearly, a weighted blowup depends on $n$ and on the choice of coordinates.
For $n=1$ the above defined map is the usual blowup of $y\in Y$.
From easy local computations (see \cite[\S 1]{Morrison-1985}) we obtain the following.

\begin{stheorem}{\bf Lemma.}\label{Lemma-Morrison}
Let $y\in Y$ be a smooth surface germ and let $f: X\to Y$  be a weighted blowup with weights $(1,n)$,
$n\ge 2$.
Let $E=f^{-1}(y)$ be the exceptional divisor and let $\pi: \tilde X\to X$ be the minimal 
resolution.
Then the exceptional locus of the composition  $\tilde X\to Y$
is  a simple normal divisor whose dual graph looks as follows:
\[
\bullet \lin \underbrace{\circ \lin \cdots \lin\circ}_{n-1}
\]
where the vertex $\bullet$ corresponds to a $(-1)$-curve \textup(the proper transform of $E$\textup)
and the vertices $\circ$ correspond to $\pi$-exceptional $(-2)$-curves.
In particular, 
$X$ has exactly one singular point and this point is Du Val of type $\A_{n-1}$.
\end{stheorem}

\begin{stheorem}{\bf Corollary.}\label{Corollary-Morrison}
In the above notation we have $K_X\cdot E=-1$. 
\end{stheorem}

\begin{stheorem}{\bf Theorem (\cite[Theorem 1.4]{Morrison-1985}).}\label{theorem-Morrison}
Let $X$ be a surface with Du Val singularities and let $f: X\to Y$ be an
extremal Mori contraction. Let $E\subset X$ be the exceptional divisor and let 
$y:=f(E)$. Then $Y$ is smooth at $y$ and 
$f$ is a weighted blowup of $y\in Y$ with weights $(1,n)$ for some $n\ge 1$.
\end{stheorem}

\end{mdefinition}

\begin{mdefinition}{\bf Contractions of surfaces with \T-singularities.}
The following is the local variant of Theorem \ref{th-1-compl} below.
 \begin{stheorem}{\bf Proposition \cite[Prop. 4.7]{Prokhorov-2004b}.} 
\label{proposition-T-contractions}
Let $X$ be a surface with \T-singularities and let $f:X\to Y$ be a contraction such that
$-K_X$ is $f$-ample. Then, near each fiber $f^{-1}(y)$, $y\in Y$, there exists 
a $1$-complement of $K_X$. 
 \end{stheorem}

\begin{stheorem}{\bf Corollary.} 
\label{corollary-T-contractions}
Let $X$ be a surface with \T-singularities and let $f:X\to Y$ be a birational contraction such that
$-K_X$ is $f$-ample. If the fiber $f^{-1}(y)$ is not a point, then $y\in Y$ is a cyclic quotient
singularity \textup(or smooth\textup).
\end{stheorem}

\begin{proof}
Let $D\in |-K_X|$ be a $1$-complement of $K_X$ near the fiber. 
If $X$ has only Du Val singularities, then $Y$ is smooth at $y$ by Theorem \ref{theorem-Morrison}.
So we assume that $K_X$ is not Cartier near $f^{-1}(y)$. Then $D\neq 0$ and $D\cap f^{-1}(y)\neq \emptyset$.
Denote $D_Y:=f_*D$. 
Since $D\sim -K_X$ is $f$-ample, we have $\Supp(D)\not\subset  f^{-1}(y)$.
Hence $D_Y\neq 0$.
The pair $(Y,D_Y)$ is lc because so $(X,D)$ is.
Moreover, $K_Y+D_Y\sim f_*(K_X+D)\sim 0$.
By the classification \ref{2-dim-lc-pairs} the point $y\in Y$ is a cyclic quotient
singularity. 
\end{proof}

\begin{sdefinition} {\bf Warning.}
In general it is not true that the singularities of $Y$ are of type \T. 
\end{sdefinition}
\end{mdefinition}

\section{$\E$- and $\D$-singularities}
\label{section-singularities}

\begin{mtheorem}{\bf Proposition.}
\label{prop-DE}
Let $X$ be a del Pezzo surface with 
at worst quotient singularities and such that $\dim |-K_X|>0$.
Then $X$ has no Du Val points of type $\D_n$ or $\E_n$
contained in $\Bs |-K_X|$.
\end{mtheorem}

\begin{mtheorem}{\bf Corollary.}\label{corollary-DE}
Let $X$ be a del Pezzo surface with \T-singularities.
Then $X$ has no Du Val points of type $\D_n$ or $\E_n$
contained in $\Bs |-K_X|$.
\end{mtheorem}

\begin{mdefinition}
Let $P\in X$ be a Du Val point of type $\D_n$ or $\E_n$
such that $P\in \Bs |-K_X|$.
Let $D\in |-K_X|$ be a general member.
Write $D=\sum d_i D_i$, where the $D_i$ are prime divisors and $d_i>0$.
\end{mdefinition}

\begin{mtheorem}{\bf Lemma.}
Notation as above.
\begin{enumerate}
\item 
For any component $D_i$ of $D$ 
we have $D_i^2\ge 0$.
\item 
If $D_i^2= 0$ for some $D_i\subset D$, then 
$P$ is the only singular point of $X$ lying on $D_i$,
the pair $(X,D_i)$ is lc, and $d_i=1$.

\item 
All the components $D_i$ pass through $P$ and do not meet 
each other elsewhere.
\end{enumerate}
\end{mtheorem}
\begin{proof}
Let $D_i$ be a component passing through $P$. Assume that $D_i^2< 0$.
Then $D_i$ generates a birational extremal ray.
If $K_X$ is not Cartier along $D_i$, then by \cite[Cor. 4.3]{Prokhorov-2004b} $X$ has no Du Val 
points of type $\D_n$ or $\E_n$ on $D_i$, a contradiction.
If $K_X$ is  Cartier along $D_i$, then $X$ has only Du Val singularities in a neighborhood 
of $D_i$ and
we have a contradiction by Theorem  \ref{theorem-Morrison} (because our ray is $K_X$-negative).
This proves (i) modulo (iii).

Now assume that $D_i^2=0$ and $D_i\ni P$.
Then $D_i$ generates a contractible $K_X$-negative extremal face.
Thus there is a contraction $f:X\to Z$, where $Z$ is a smooth curve, such that $D_i=f^{-1}(z)_{\red}$
for some $z\in Z$. 
Since $D\in |-K_X|$ is a general member, 
the scheme fiber $f^*z$ is not contained in $D$. So, $f^*z\neq D_i$,
i.e. $f^*z$ is not reduced.
By \cite[Cor. 4.3]{Prokhorov-2004b}\ $K_X$ is Cartier along $D_i$.
Since $-K_X\cdot f^*z=2$, we have $-K_X\cdot D_i=1$ and $D_i$ is
a fiber of multiplicity $2$. Since $2D_i =f^*z$ is not contained in $D$, we have $d_i=1$.
Finally, by \cite[Prop. 7.1.3, Th. 7.1.12]{Prokhorov-2001} the pair $(X,D_i)$ is lc.
This proves (ii) modulo (iii).

Assume that (iii) does not hold. Then
there is a component $D_i$ such that 
$P\in D_i$ and $D_i\cap D_j\ni Q\neq P$
for some $j\neq i$ (because $\Supp D$ is connected). Put $D':=D-d_iD_i$.
By the above, $D_i^2\ge 0$. 
Hence the divisor $-(K_X+D_i+D')\sim (d_i-1)D_i$ is nef. Then by 
the adjunction we have
\[
\deg  \Diff_{D_i}(D')\le - \deg K_{D_i}\le 2.
\]
On the other hand,
$\Diff_{D_i}(D')\ge P+Q$. 
Hence, $\Diff_{D_i}(D')= P+Q$ and $(X,D_i+D')$ is lc near $D_i$
\cite[17.6]{Utah}. Since $P\in X$ is not a cyclic quotient
singularity,  the pair $(X,D_i)$ is strictly lc (i.e. lc but not plt) 
at $P$ (see \ref{2-dim-lc-pairs}).
In particular,
 no component of $D'$ pass through $P$.
Hence, $K_X+D_i+D'$ is not Cartier at $P$ (see \ref{2-dim-lc-pairs-lc-D}),
so $D\neq D_i+D'$ and $d_i>1$.
In this case, $D_i^2>0$ by (ii).
Then 
\[
(K_X+D_i+D')\cdot D_i=-(d_i-1)D_i^2<0
\]
and $\deg \Diff_{D_i}(D')<2$. The contradiction proves (iii). 
\end{proof}

\begin{proof}[Proof of Proposition {\rm \ref{prop-DE}}]
Let now $Q\in X$ be a non-Du Val point and let $D_i$ be a component
of $D$ passing through $Q$. Write $D=d_iD_i+D'$.

If  $D=D_i$, then
by \ref{2-dim-lc-pairs-plt} the pair $(X,D)$ is not plt at $Q$.
Hence, as above, $\Diff_{D}(0)= P+Q$.
By \cite[17.6]{Utah} $(X,D)$ is lc near $D$.
But then $K_X+D$ is not Cartier near $Q$, a contradiction.

Therefore, $D\neq D_i$.
Since $\dim |D|>0$, $D'\neq 0$ (and $D'$ has a reduced movable component). 
Note that the divisor $-(K_X+D_i+D')\sim (d_i-1)D_i$ is nef. Then by 
the adjunction we have
\[
\deg  \Diff_{D_i}(D')\le - \deg K_{D_i}\le 2.
\]
Since the coefficients of $\Diff_{D_i}(0)$ are $\ge 1/2$
and $(X,D_i)$ is not plt at $P$, $\Supp \Diff_{D_i}(D')=\{P,\, Q\}$.
Write $\Diff_{D_i}(0)=a_0P+bQ$ and $\Diff_{D_i}(D')=aP+bQ$.
Since $(X,D_i)$ is not plt at $P$, $a>a_0\ge 1$.
Hence $b<1$ and $(X,D_i)$ is plt at $Q$.
Thus, $b=1-1/m$ for some $m\ge 3$ (because 
$Q\in X$ is not Du Val of type $\A_1$)
and $a\le 4/3$.
If $(X,D_i)$ is not lc at $P$, then $a_0\ge 1+1/l$,
where $l$ is the minimal positive integer such that 
$l(K_X+D_i)$ is Cartier at $P$.
Recall that 
for the Weil divisor class group of a Du Val singularity $(X,P)$
we have 
\[
\begin{array}{c||ccccc}
(X,P)&
\D_{2n+1}&\D_{2n}&\E_6&\E_7&\E_8
\\[5pt]
\hline
\\[-5pt]
\operatorname{Cl}(X,P)&\ZZ/4\ZZ &\ZZ/2\ZZ\oplus \ZZ/2\ZZ& \ZZ/3\ZZ& \ZZ/2\ZZ&0
\end{array}
\]
(see, e.g., \cite{Brieskorn-1967-1968}).
So in our case we have $a_0\ge 5/4$ and 
$a\ge 5/4+1/4=3/2$, a contradiction. 

Thus we may assume that $(X,D_i)$ is  lc at $P$.
In particular, $(X,P)$ is of type $\D_n$.
Then $a_0=1$. Since $a\le 4/3$, we have only one possibility:
$b=2/3$, $a=5/4$, and $2D'$ is not Cartier at $P$. 
Moreover, $D'$ is irreducible (and reduced),
$(X,D')$ is not lc at $P$ (see \ref{2-dim-lc-pairs-lc-D}),
and so $\Diff_{D'}(0)\ge \frac54P$. 
Again by \ref{2-dim-lc-pairs-lc-D} $2D_i$ is Cartier at $P$.
Hence, 
\[
\Diff_{D'}(D_i)\ge \left(\frac54+\frac{d_i}2\right)P>2P,
\]
a contradiction.
\end{proof}

\section{Existence of $1$-complements}
\label{section-1-complements}

In this section we prove the following important fact (cf. \cite[Th. 7.1]{Hacking-Prokhorov-2010}).
\begin{mtheorem}{\bf Theorem.}
\label{th-1-compl}
Let $X$ be a del Pezzo surface with \T-singularities. 
Then there exists a $1$-complement of $K_X$. 
\end{mtheorem}

We need a few preliminary facts.

\begin{mdefinition}{\bf Definition \cite[\S 2]{Prokhorov-Shokurov-2009}.}
\label{CY->F}
Let $X$ be a normal projective variety. We say that
$X$ is \emph{FT} \textup(\emph{Fano type}\textup) if
there is a $\QQ$-boundary $\Delta$ such that 
$(X,\Delta)$ is a klt (Kawamata log terminal)  log Fano.
\end{mdefinition}

\begin{stheorem}{\bf Proposition \cite[\S 2]{Prokhorov-Shokurov-2009}.}
\label{CY-MMP}
Let $X$ be an FT variety. 
\begin{enumerate}
\item 
The Mori cone $\NE(X)$ is polyhedral and has contractible faces.
\item 
If $f: X\to Z$ be any contraction of normal varieties.
Then $Z$ is FT. In particular, the FT property is preserved under 
MMP.
\item
Let $\Xi$ be a boundary on $X$ such that $(X,\Xi)$ is lc and 
$-(K_X+\Xi)$ is nef.
Let $f: Y\to X$ be a birational extraction such that
$a(E,X,\Xi)< 0$ for every $f$-exceptional divisor $E$.
Then $Y$ is also FT.
\item 
Assume the LMMP in dimension $\dim X$.
Then the $D$-MMP works on $X$ with respect to any
divisor $D$.
\end{enumerate}
\end{stheorem}

\begin{mtheorem}{\bf Proposition.}
\label{prop-1-compl}
Let $(Y,C)$ be a log pair where $Y$ is an FT surface
and $C$ is an irreducible curve.
Assume that $(Y,C)$ is plt, $-(K_Y+C)$ is nef and big,
and $|-(K_Y+C)|\neq \emptyset$. Then one of the following holds:
\begin{enumerate}
\item 
$K_Y+C$ has a $1$-complement,
\item
$Y$ has three or four singular points on $C$ and either
\begin{enumerate}
\item 
$C^2<0$, $K_Y\cdot C\ge 0$, or
\item 
$\dim |-K_Y|=0$ and $-K_Y\sim bC$, $b\ge 2$.
\end{enumerate}
\end{enumerate}
\end{mtheorem}
\begin{proof}
First of all note that the curve $C$ is smooth 
(see \ref{2-dim-lc-pairs-plt}).
By Proposition \ref{prodolj} we can extend complements from 
$(C,\Diff_{C}(0))$ to $Y$.
Thus, for (i), it is sufficient to show existence of a $1$-complement of $K_{ C}+\Diff_{C}(0)$. 
Assume the converse and write
\[
\Diff_C(0)=\sum \left(1-\frac 1{m_i} \right)P_i, \quad
\deg K_C+\deg \Diff_C(0)=(K_X+C)\cdot C\le 0
\]
(see \ref{2-dim-lc-pairs-plt}). 
Thus, $\sum (1-1/m_i)\le 2$.
Since, by our assumption, the log divisor $K_{ C}+\Diff_{C}(0)$ is not $1$-complementary,
easy computations \cite[19.5]{Utah} 
show that $\Diff_{C}(0)$ is supported in 
three or four points $P_i$.
In particular, $\deg K_C<0$ and so $C\simeq \PP^1$.

Assume that $C^2\le 0$. Then $C$ generates an extremal face.
Since $Y$ is FT, this extremal face is contractible:
there is a contraction $\varphi: Y\to Y'$ such that $y:=\varphi(C)$ 
is a point. By Lemma \ref{lemma-2-points} below $K_Y\cdot C\ge 0$.
If $C^2=0$, then $Y'$ is a curve, $\varphi$ is a rational curve fibration, 
and $C=\varphi^{-1}(y)_{\red}$. 
In this case, $K_Y\cdot C<0$, a contradiction. 
Thus $C^2<0$ and we are 
in the case (iia).

Assume that $C^2>0$.
Let $D\in |-(K_Y+C)|$ be a general member. Write $D=aC+D'$,
where $a\ge 0$ and $C$ is not a component of $D'$.
If $D'=0$, we get case (iib). (Here $b=a+1\ge 2$ because
$K_X+C$ is not Cartier near singular points on $C$, see  \ref{2-dim-lc-pairs-plt}.) 
Thus we may assume that $D'\neq 0$.
Since the support of 
$$
(a+1)C+D'\in |-K_Y|
$$ 
is connected, $D'$ meets $C$. 
Further,
\begin{equation}
\label{eq-diff}
\deg \Diff_C( D')=-\deg K_C + (K_Y+ C+ D')\cdot C=2-(a-1)C^2\le 2. 
\end{equation}
By the above, $\Diff_C(D')$ has 
at least one point of multiplicity $\ge 1$
(and multiplicities of all points are $\ge 1/2$).
Since $\Diff_C(D')\ge \Diff_C(0)$, 
the only possibility is 
\[
\Diff_C(D')=P_1+\frac12 P_2+\frac12 P_3,
\]
where $P_1\in C\cap \Supp(D')$ and $P_2,\, P_3 \notin C\cap \Supp(D')$.
By \ref{2-dim-lc-pairs-plt} $K_Y+C+D'$ is not Cartier at $P_2$ and $P_3$.
Thus $K_Y+C+D'\not \sim 0$ and $a>1$.
On the other hand, $\deg \Diff_C(D')=2$, so by \eqref{eq-diff}
$a=1$, a contradiction.
\end{proof}

\begin{mtheorem}{\bf Lemma \cite[Prop. 7.1.12]{Prokhorov-2001}.}
\label{lemma-2-points}
Let $S\to Z$ be a $K$-negative extremal contraction
from a surface $S$ with log terminal singularities, 
where $Z$ is not a point.
Then $S$ has at most two singular points on each fiber.
\end{mtheorem}

\begin{proof}[Proof of Theorem {\rm \ref{th-1-compl}}]
Let $X$ be a del Pezzo surface with 
at worst quotient singularities and such that $\dim |-K_X|>0$
and let $D\in |-K_X|$ be a general member.
Take $t\in \QQ$ so that $(X,tD)$ is maximally lc.
If $t=1$, then $K_X+D$ is a $1$-complement.
So from now on we assume that $t<1$.

\begin{mdefinition}
Consider the case where $(X,tD)$ is plt. 
Write $tD=C+B$, where $C:=\down{tD}\neq \emptyset$
and $B$ is an effective fractional divisor.
Since $X$ is an FT variety,
we can run $-(K+C)$-MMP and obtain
\[
\varphi: X \longrightarrow \bar X.
\]
Since 
\[
-(K_X+C)\equiv B-(1-t)K_X, 
\]
all the contractions are $B$-negative.
Hence they are birational and we end up with a model 
$(\bar X,\bar C)$ such that $-(K_{\bar X}+\bar C)$ is nef.
We have 
\[
-(K_{\bar X}+\bar C)\equiv \bar B-(1-t)K_{\bar X}, 
\]
where
$-K_{\bar X}$ is ample and $\bar B:=\varphi_*B$ is effective.
Hence the divisor 
$-(K_{\bar X}+\bar C)$ is big. Further, 
\[
K_X+C+B\equiv -(1-t)D\equiv (1-t)K_X. 
\]
Hence all the contractions 
in $\varphi$ are $(K+C+B)$-negative.
Therefore, $(\bar X, \bar C+\bar B)$ is plt and so is 
$(\bar X, \bar C)$. 
So, $\bar B\neq 0$ and $\bar D:=\varphi_*D\neq \bar C$.
Apply Proposition \ref{prop-1-compl} to $(\bar X,\bar C)$.
The case (iia) does not occur because 
$-K_{\bar X}$ is ample and the case (iib) does not occur because
\[
\dim |-K_{\bar X}|\ge \dim |-K_X|>0.
\]
Hence, there exists a $1$-complement of $K_{\bar X}+\bar C$. 
By 
Proposition \ref{bir-prop} we can pull back $1$-complements 
from $\bar X$ to $X$.
\end{mdefinition}

\begin{mdefinition}
Now consider the case where $(X,tD)$ is not plt.
Put $B:=tD$.
Consider an inductive plt blowup \cite[Prop. 3.1.4]{Prokhorov-2001}\quad
$\delta: \hat X \to X$, that is, a birational extraction such that 
$\rho(\hat X/X)=1$ and 
\[
K_{\hat X}+\hat B+C=\delta^*(K_X+B),
\]
where $C$ is the (irreducible) exceptional divisor
and $\hat B$ is the strict transform of $B$.
Moreover, the pair $(\hat X, C +(1-\epsilon) \hat B)$ is plt 
for any $\epsilon >0$. 
Write 
\[
K_{\hat X}+\hat D+aC=\delta^*(K_X+D), 
\]
where $\hat D$ is 
the strict transform of $D$ and $a> 1$.
Then $\hat D+aC\in |-K_{\hat X}|$, so
$\dim |-K_{\hat X}|>0$. By Proposition \ref{CY-MMP} 
the variety $\hat X$ is FT. Run the $-(K+C)$-MMP.
As above all the contractions are $\hat B$-negative.
So we end up with a model $(\bar X,\bar C)$ where
$-(K_{\bar X}+\bar C)$ is nef and big (and $\bar C\neq 0$):
\[  
\xymatrix{
&\hat X\ar[dr]^{\varphi}\ar[dl]_{\delta}&
\\
X&&\bar X
}
\]
(the case $\hat X=\bar X$ is not excluded).
Since $\NE(\hat X)$ is polyhedral,
$-(K_{\hat X}+ C +(1-\epsilon) \hat B)$ is ample 
for some $0<\epsilon \ll 1$. Hence the plt property of 
the pair $(\hat X, C +(1-\epsilon) \hat B)$ is preserved.
In particular, $(\bar X,\bar C)$ is plt.
Apply Proposition \ref{prop-1-compl} to $(\bar X,\bar C)$.
The case (iib) does not occur because
\[
\dim |-K_{\bar X}|\ge \dim |-K_{\hat X}|>0.
\]
Assume that we are in the case (iia). 
Then $\bar C^2<0$ and $K_{\bar X}\cdot \bar C\ge 0$.
In particular, $\bar C$ is contractible: there is a contraction 
$\psi : \bar X\to \check X$ of $\bar C$, where $\check X$
is an FT surface. 
As in \cite[Proof of Th. 7.1]{Hacking-Prokhorov-2010} we see that
$\check P:=\psi(\bar C)\in \check X$ is a singular point and 
it is not a cyclic quotient singularity.
According to Zariski's main theorem the composition 
$\upsilon=\psi\comp\varphi\comp \delta^{-1}: X\dashrightarrow 
\check X$ is a morphism.
By Corollary \ref{corollary-DE} $\upsilon$ is not an isomorphism
(because $\delta(C)\in \Bs |-K_X|$).
Since $X$ is a del Pezzo, $\upsilon$ is 
a $K$-negative contraction.
On the other hand, by Corollary \ref{corollary-T-contractions}
$\check P$ is a cyclic quotient singularity,
a contradiction.
Therefore, the case (iia) does not occur 
and so there exists a $1$-complement of $K_{\bar X}+\bar C$. 
Now as above by 
Proposition \ref{bir-prop}  we can pull back $1$-complements 
from $\bar X$ to $X$.
\end{mdefinition}
\end{proof}

\section{Tori actions}
\label{section-Tori}
For a normal projective surface $X$ we denote by $\varrho(X)$ the \textit{numerical Picard number}, that is,
the rank of the group of Weil divisors modulo numerical equivalence.
Clearly, $\varrho(X)\ge \rho(X)$ 
and the equality holds if $X$ is $\QQ$-factorial.
For a $\QQ$-divisor $D=\sum d_iD_i$ on $X$ we denote
\begin{equation*}
\begin{array}{lll}
\| D \|&:=&\sum d_i,
\\[7pt]
\varsigma(X, D)&:= &\varrho(X)+2-\|D\|.
\end{array}
\end{equation*}
We say that a log pair $(X,D)$ is \textit{toric} if $X$ is a toric variety and 
$D$ is the (reduced) invariant boundary. 
We say that a log pair $(X,D)$ admits an \textit{effective $\CC^*$-action} if 
the variety $X$ admits such an action so that 
the divisor $D$ is $\CC^*$-invariant.

The statements (i) and (ii) of the following theorem
ware proved by Shokurov in much more general form \cite{Shokurov-2000}.
For the convenience of the reader we provide simplified complete proofs.

\begin{mtheorem}{\bf Theorem.}
\label{theorem-toric}
Let $(X, D)$ be a projective normal log surface 
such that $D$ is an integral \textup(effective\textup)
divisor, the pair $(X, D)$ is lc, and $K_X+D \sim 0$. Then
\begin{enumerate}
 \item 
$\varsigma(X, D)\ge 0$;
 \item 
if the equality holds, then $(X,D)$ is toric; 
 \item 
if $\varsigma(X, D)=1$, then $(X,D)$ admits an effective $\CC^*$-action. 
\end{enumerate}
\end{mtheorem}

\begin{sdefinition}{\bf Remark.}
Let $X$ be a projective normal surface and let $D\in |-K_X|$ be a divisor such that 
the pair $(X,D)$ is lc. Then
the property $\varsigma(X, D)= 0$ characterizes toric pairs.
On the other hand, the condition $\varsigma(X, D)\le  1$ is sufficient 
but not necessary 
for $(X,D)$ to admit an effective $\CC^*$-action.
For example, the product $C\times \PP^1$, where $C$ is an elliptic curve, 
admits an effective $\CC^*$-action
but for any $D\in |-K_X|$
we have $\varsigma(X, D)\ge 2$.
\end{sdefinition}

The rest of this section is devoted to the proof of Theorem
\ref{theorem-toric}.
We will use the following fact which 
is an easy consequence of the definition.

\begin{mtheorem}{\bf Lemma.}
\label{lemma-sigma-extremal}
Let $\varphi: Y'\to Y$ be a birational morphism of normal surfaces,
 and let
$D'$ be a reduced boundary on $Y'$. 
Denote by $N(\varphi, D')$ the number of $\varphi$-exceptional
curves that are not contained in the support of $D'$. 
Then
\[
\varsigma(Y',D')=
\varsigma(Y,g_*D')+N(\varphi, D').
\]
\end{mtheorem}

\begin{mdefinition}
Let $(X, D)$ be a projective log surface such that $(X, D)$ is lc, $K_X+D \sim 0$, and
\begin{equation}
\label{eq-main-contr}
\varsigma:=\varsigma(X, D)\le 1.
\end{equation}
Let $f: (X',D') \to (X,D)$ be a 
minimal dlt modification, that is, a birational map such that 
the log pair $(X',D')$ is dlt (divisorial log terminal \cite[2.13]{Utah}),
\[
K_{X'}+D'\sim f^* (K_X+D),\qquad f_*D'=D, 
\]
and any $f$-exceptional divisor 
has multiplicity $1$ in $D'$ (see e.g. \cite[Prop. 21.6.1]{Utah},
\cite[Prop. 3.1.2]{Prokhorov-2001}). By Lemma \ref{lemma-sigma-extremal}
\[
\varsigma(X', D')=\varsigma(X, D)=\varsigma\le 1.
\]
Hence, \eqref{eq-main-contr} holds for $(X',D')$.
Since $(X',D')$ is dlt, $X'$ is non-singular near $D'$ \cite[Prop. 16.6]{Utah}.
Moreover, $X'$ has at worst Du Val singularities outside of $D'$.

\begin{sdefinition}\label{MMP}
Run the $K$-MMP:
\[
(X',D')=(X^{(1)}, D^{(1)}) \overset{\varphi_1}\longrightarrow (X^{(2)}, D^{(2)}) \overset{\varphi_2} \longrightarrow \cdots 
\overset{\varphi_{l-1}}\longrightarrow (X^{(l)}, D^{(l)})= (Y, D_Y). 
\]
Let $E^{(i)}\subset X^{(i)}$ be the $\varphi_i$-exceptional
divisor. 
\end{sdefinition}
\end{mdefinition}

\begin{mtheorem}{\bf Claim.}\label{claim-Morrison}
For each $i=1,\dots,l$ we have 
\begin{enumerate}
\item 
$X^{(i)}$ has at worst Du Val singularities;
\item
$X^{(i)}$ is non-singular near $D^{(i)}$;
\item
$\varphi_{i}$ is the weighted blowup 
with weights $(1,n)$,  $n\ge 1$ of a smooth point $\varphi_{i}(E^{(i)}) \in X^{(i+1)}$;
\item 
$D^{(i)}$ is a simple normal crossing divisor;
\item
$E^{(i)}\cdot D^{(i)}=1$.
\end{enumerate}
\end{mtheorem}

\begin{proof}
One can prove (i)-(iii) by induction on $i$ using the following scheme:
\[
\text{ (i)${}_{i}$, (ii)${}_{i}$  $\Longrightarrow$  (iii)${}_{i}$, (i)${}_{i+1}$, (ii)${}_{i+1}$.}
\]
Indeed, if (i)${}_i$ holds, then $\varphi_i$ is a weighted blowup by Theorem \ref{theorem-Morrison}
and so $X^{(i+1)}$ is smooth at $\varphi_i(E^{(i)})$.

Since the pair $(X^{(i)}, D^{(i)})$ is lc, (ii) implies (iv) and
(v) follows from Corollary \ref{Corollary-Morrison}
because $ D^{(i)}\sim -K_{X^{(i)}}$.
\end{proof}

\begin{mtheorem}{\bf Claim.}
For each $i=1,\dots,l$ we have
$\varsigma(X^{(i)}, D^{(i)})\le \varsigma\le 1$,
in particular, $D^{(i)}\neq 0$.
\end{mtheorem}
\begin{proof}
It follows from Lemma \xref{lemma-sigma-extremal}. 
\end{proof}

By Lemma \xref{lemma-sigma-extremal} and because $X^{(i)}$ is non-singular near $D^{(i)}$,
on each step we have one of the following possibilities:
\begin{sdefinition}\label{claim-MMP-2-possibilities-i}
$\varsigma(X^{(i+1)}, D^{(i+1)})=\varsigma(X^{(i)}, D^{(i)})$,
$E^{(i)}\subset D^{(i)}$, 
and $\varphi_i$ is the usual blowup of a singular point of $D^{(i)}$;
\end{sdefinition}
\begin{sdefinition}\label{claim-MMP-2-possibilities-ii}
$\varsigma(X^{(i+1)}, D^{(i+1)})=\varsigma(X^{(i)}, D^{(i)})-1$, and
$E^{(i)}\not\subset D^{(i)}$.
\end{sdefinition}

\begin{stheorem}{\bf Corollary.}\label{corollary-MMP-2-possibilities}
Suppose that we are in the case \xref{claim-MMP-2-possibilities-i} above.
Furthermore suppose that $X^{(i+1)}$ admit an action of a 
connected algebraic group $G$ so that the boundary $D^{(i+1)}$
is $G$-invariant. Then the action lifts to $X^{(i)}$ so that $D^{(i)}$
is $G$-invariant.
\end{stheorem}

\begin{stheorem}{\bf Corollary.}\label{corollary-MMP-2-possibilities-a}
Suppose that we are in the case \xref{claim-MMP-2-possibilities-ii} above.
Furthermore suppose that $(X^{(i+1)},D^{(i+1)})$ 
is a toric surface.
Then the action of some one-dimensional 
subtorus $\mathcal T$ lifts to 
 $X^{(i)}$ so that $D^{(i)}$
is $\mathcal T$-invariant.
\end{stheorem}

\begin{proof}
Since $E^{(i)}\cdot D^{(i)}=1$ and $E^{(i)}\not\subset D^{(i)}$, the curve $E^{(i)}$ meets 
only one component $D^{(i)}_0\subset D^{(i)}$ so that $E^{(i)}\cdot D^{(i)}_0=1$.
Let $\pi: \bar X^{(i)}\to X^{(i)}$ be the minimal resolution near $E^{(i)}$.
By Lemma \ref{Lemma-Morrison}
the dual graph of $\bar X^{(i)}$ has the following form:
\[
\Box \lin \bullet \lin \underbrace{\circ \lin \cdots \lin\circ}_{n-1}
\]
where $\bullet$ corresponds to $E^{(i)}$, $\Box$ corresponds to $D^{(i)}_0$, 
and the vertices $\circ$ correspond to $\pi$-exceptional $(-2)$-curves.
Thus $\bar X^{(i)}$ is obtained from $X^{(i+1)}$ by making successive blowups of a fixed point
on the proper transform of $D^{(i)}_0$.
The stabilizer of this point is a one-dimensional subtorus in the big torus acting on $X^{(i+1)}$.
\end{proof}

\begin{mdefinition}
At the end of our MMP \ref{MMP} we get a log surface $(Y,D_Y)$
admitting a fiber type extremal $D_Y$-positive contraction $h: Y\to Z$. 
Moreover,
\begin{equation}
\label{eq-main-contr-Y}
\varsigma(Y, D_Y)\le \varsigma\le 1.
\end{equation}
Recall that $Y$ is non-singular near $D_Y$.
In particular, all the component of $D_Y$ are Cartier divisors.
Moreover, $K_Y+D_Y\sim 0$ and $Y$ has at worst Du Val singularities outside of $D_Y$.
\end{mdefinition}

\begin{mdefinition}\label{toric-del_Pezzo-case}
First we consider the case where $Z$ is a point.
Then $Y$ is a del Pezzo surface with at worst Du Val singularities and 
$\Pic (Y)\simeq \ZZ$. 
In particular, $\varrho(Y, D_Y)=1$.
Since $\|D_Y\|\ge 2$, the divisor $-K_Y$ is not a primitive element
of $\Pic(Y)$. In this case, $Y$ is either a projective plane $\PP^2$ or 
a singular quadric $\PP(1,1,2)$ (see e.g. \cite[Lemma 6]{Miyanishi-Zhang-1988}). Moreover
 we have one of the following:
\begin{sdefinition}
\label{end-3-components}
$Y\simeq\PP^2$, $D_Y=D_1+D_2+D_3$, where $D_i$ are lines in general position,
$\varsigma(Y, D_Y)=0$;
\end{sdefinition}
\begin{sdefinition}
\label{end-2-components-P2}
$Y\simeq\PP^2$, $D_Y=D_1+D_2$, where $D_1$ is a line and $D_2$ is a conic meeting 
$D_1$ transversely at two distinct points,
$\varsigma(Y, D_Y)=1$; 
\end{sdefinition}
\begin{sdefinition}
\label{end-3-components-P112}
$Y\simeq\PP(1,1,2)$, $D_Y=D_1+D_2$, where the class of $D_i$
generates $\Pic(Y)$ and again $D_1$ and $D_2$ meet each other transversely at two distinct points,
$\varsigma(Y, D_Y)=1$. 
\end{sdefinition}

\begin{claim*}
The pair $(Y, D_Y)$ is toric in the case \xref{end-3-components}
and admits an effective $\CC^*$-action in cases
\xref{end-2-components-P2} and \xref{end-3-components-P112}.
\end{claim*}

\begin{proof}
Modulo change of coordinates $x$, $y$, $z$  in $\PP^2$ or $\PP(1,1,2)$   we have 
\begin{itemize} 
\item[\ref{end-3-components}] $\Longrightarrow$ $D_Y=\{xyz=0\}$,
\item[\ref{end-2-components-P2}] $\Longrightarrow$ $D_Y=\{(xy-z^2)z=0\}$,
\item[\ref{end-3-components-P112}] $\Longrightarrow$  $D_Y=\{(xy-z)z=0\}$.
\end{itemize}
Then the statement of the claim is an easy exercise.
\end{proof}

In all cases \ref{end-3-components}--\ref{end-3-components-P112}
we have $\varsigma(Y, D_Y)\ge 0$.
This proves (i) of Theorem \ref{theorem-toric}. 
Moreover, if $\varsigma(X, D)= 0$,
then 
we are in the case \ref{end-3-components}.
In particular, $(Y,D_Y)$ is a toric surface.
Since $\varsigma(Y, D_Y)=\varsigma(X', D')=0$,
on each step of our MMP we have the
possibility \ref{claim-MMP-2-possibilities-i}.
By Corollary \ref{corollary-MMP-2-possibilities} both 
$(X',D')$ and $(X,D)$ are toric.

Now assume that $\varsigma(X, D)= 1$.
If moreover $\varsigma(Y, D_Y)=1$, then $(Y,D_Y)$ is 
of type \ref{end-2-components-P2} or \ref{end-3-components-P112}
and each step of our MMP is of type \ref{claim-MMP-2-possibilities-i}.
By Corollary \ref{corollary-MMP-2-possibilities} 
the action of the corresponding one-dimensional torus 
lifts to $X'$. Hence $(X',D')$ and $(X,D)$ admit effective $\CC^*$-actions.
Finally assume that $\varsigma(X, D)= 1$ and $\varsigma(Y, D_Y)=0$.
Then $(Y,D_Y)$ is toric and all but one steps of our MMP are of type 
\ref{claim-MMP-2-possibilities-i}.
As above we can apply Corollaries \ref{corollary-MMP-2-possibilities} 
and \ref{corollary-MMP-2-possibilities-a} to conclude that 
$(X,D)$ admits an effective $\CC^*$-action.
\end{mdefinition}

\begin{mdefinition}
Now consider the case where 
$Z$ is a curve. Then $Z$ is smooth and $h: Y\to Z$ is a rational curve fibration
with $\Pic(Y/Z)\simeq \ZZ$.
For a general fiber $F$ we have 
\[
D_Y\cdot F=-K_Y\cdot F=2. 
\]
Let $D_0$ be a $h$-horizontal component of $D_Y$. We claim that $D_0$ is a section.
Indeed, assume that $D_0$ is a double section. Then by the adjunction formula
\[
{D_0}\cdot (D_Y-D_0)=-{D_0}\cdot (K_Y+D_0)= -\deg K_{D_0}\le 2.
\]
Since $\|D_Y\|\ge 3$ and $D_Y$ is a simple normal crossing divisor, it 
has at least two 
vertical components $D_i$ with $D_0\cdot D_i=2$. Thus ${D_0}\cdot (D_Y-D_0)\ge 4$,
a contradiction. 

Hence, $D_0$ is a section. 
Then $D_Y$ has another $h$-vertical component $D_1$ which is also a 
section of $h$.
Since $D_0$ is a Cartier divisor, 
$h: Y\to Z$ is a smooth $\PP^1$-fibration.
If $D_0$ is not a rational curve, then as above by adjunction 
$\deg K_{D_1}= 0$ and $D_0$
is disjoint from $D_Y-D_0$. On the other hand, 
$D_Y-D_0$ has at least one $h$-vertical component, a contradiction. 
Hence, $D_0$ is a smooth rational curve and $Y$ is a Hirzebruch surface
$\FF_e$, $e\neq 1$. Let $\Sigma$ be the minimal section of $\FF_e$ and let 
$F$ be a fiber.
Since $D_i\cdot (D_Y-D_i)= -\deg K_{D_i}=2$ for each component 
$D_i\subset D_Y$, we have one of the following possibilities
(up to permutation of $D_0$ and $D_1$):

\begin{sdefinition}
\label{end-ruled-4-components}
$D_0\cdot D_1=0$, $\|D_Y\|=4$, $D_0=\Sigma$, $D_1\sim \Sigma+eF$,
$D_2$ and $D_3$ are distinct fibers, $\varsigma(Y, D_Y)=0$;
\end{sdefinition}

\begin{sdefinition}
\label{end-ruled-3-components}
$D_0\cdot D_1=1$, $\|D_Y\|=3$, $D_0=\Sigma$, $D_1\sim \Sigma+(e+1)F$,
$D_2$ is a fiber, $\varsigma(Y, D_Y)=1$.
\end{sdefinition}

Then we can complete the proof similar to \ref{toric-del_Pezzo-case}
by using the following.
\end{mdefinition}

\begin{claim*}
The pair $(Y, D_Y)$ is toric in the case \xref{end-ruled-4-components}
and admits an effective $\CC^*$-action in the case
\xref{end-ruled-3-components}.
\end{claim*}
\begin{proof}
The statement is obvious in the case $e=0$, so we assume that $e\ge 2$.
Let $\pi: (Y,D_Y)\to (Y',D_Y')$ be the contraction of the negative section.
Then $Y'$ is the weighted projective plane $\PP(1,1,e)$. We may assume that in suitable orbifold coordinates 
$x,\, y,\, z$ the boundary
$D_Y'$ is given by the equation 
$xyf_{e}(x,y,z)=0$ (resp. $xf_{e+1}(x,y,z)=0$) in the case \xref{end-ruled-4-components}
(resp. 
\xref{end-ruled-3-components}), where $f_{d}(x,y,z)$ denotes some polynomial of weighted degree $d$.

In the case \xref{end-ruled-4-components},
since $D_Y'$ is a simple normal crossing divisor outside of the origin $(0:0:1)$
and $D_1'$ does not pass through $(0:0:1)$,
the polynomial $f_{e}(x,y,z)$ contains $z$.
Then by a coordinate change we get $f_{e}(x,y,z)=z$. Hence $(Y',D_Y')$ is toric.
Since $\pi$ is the minimal resolution, the torus action lifts to $Y$.

Similarly, in the case \xref{end-ruled-3-components} 
$f_{e+1}(x,y,z)$ contains $zy$ (because $(Y',D_Y')$ is lc).
By a coordinate change we get $f_{e+1}(x,y,z)=zy+x^{e+1}$.
Then $(Y',D_Y')$ admits an $\CC^*$-action 
$(x,y,z)\longmapsto(x,\lambda y, \lambda^{-1}z)$.
\end{proof}

\section{Proof of main theorems}
\label{section-main-theorem}
Now Theorem \ref{main} is a consequence of the following.
\begin{mtheorem}{\bf Proposition.} \label{proposition-e-main}
Let $X$ be a projective normal surface
and let $\s(X)$ be the number of its points where $K_X$ is not Cartier.
Assume that $X$ has a $1$-complement $D\in |-K_X|$. Then 
\begin{enumerate}
\item 
$\s(X)\le \varrho(X)+2$,
\item 
if $\s(X)=\varrho(X)+2$, then $X$ is toric,
\item 
if $\s(X)=\varrho(X)+1$, then $X$ admits an effective $\CC^*$-action.
\end{enumerate}
\end{mtheorem}
\begin{proof}
By the classification of log canonical singularities of pairs \cite[Thm.~4.15]{Kollar-Mori-19988}, 
$D$ is a nodal curve,
and, at each singularity $P \in X$, either $D=0$ and $P \in X$ is a Gorenstein log canonical singularity, 
or the pair $P \in (X, D)$ is locally analytically isomorphic to the pair $(\frac{1}{n}(1,a),(uv=0))$ for some $n$ and $a$.
Moreover $D$ has arithmetic genus $1$ because $2p_a(D)-2=(K_X+D) \cdot D = 0$ (note that the adjunction formula holds
because $K_X+D$ is Cartier \cite[16.4.3]{Utah}). 
Thus $D$ is either a smooth elliptic curve, or a
rational curve with a node, or a cycle of smooth rational curves.

Let $\s'$ be the number of singular points of $X$ lying on $D$.
Then 
\[
\# \Sing(D) \ge \s'\ge \s(X).
\] 
By the above $\# \Sing(D)=\|D\|$. 
Then the assertion follows from Theorem \ref{theorem-toric}.
\end{proof}

The proof of Theorem \ref{thm-ge} is essentially the same as the proof of  \cite[Theorem 1.3]{Hacking-Prokhorov-2010}.

\section{Examples}
\label{section-Examples}
A natural way to produce examples of 
del Pezzo surfaces as in (iii) 
of Theorem \ref{main} is to use deformations:

\begin{mtheorem}{\bf Theorem \cite[Prop. 3.1]{Hacking-Prokhorov-2010}.} \label{unobs}
Let $X$ be a projective surface such that $X$ has only \T-singularities and 
$-K_X$ is nef and big. 
Then there are no local-to-global obstructions to deformations of $X$.
\end{mtheorem}

Thus we can start with some known examples and construct new ones by deforming 
their singularities. The behavior of the Picard number is described by 
Noether's formula \ref{noether} and by the following

\begin{mtheorem}{\bf Proposition \cite[Prop. 2.3]{Hacking-Prokhorov-2010}.} 
\label{Tdeformation}
Let $(P \in \mathcal X)/(0 \in S)$ be a $\QQ$-Gorenstein deformation 
of a \T-singularity $P\in X$ of type $\frac{1}{dn^2}(1,dna-1)$
and let $P_1,\dots,P_l$ be all the singular points of a fiber $\mathcal X_s$, $s\in S$.
Then the possible types of $P_1,\dots,P_l\in \mathcal X_s$  are as follows:
\begin{enumerate}
 \item[a)]
$A_{d_1-1},\ldots,A_{d_l-1}$ or 
 \item[b)] 
$\frac{1}{d_1n^2}(1,d_1na-1)$, 
$A_{d_2-1},\ldots, A_{d_l-1}$,
\end{enumerate}
where $d_1,\ldots,d_l$ is a partition of $d$. 
\end{mtheorem}

\begin{sdefinition}{\bf Remark.}
In the above situation, 
the case $\Sing(\mathcal X_s)=\emptyset$ is not excluded.
This is possible only if $d=1$ and 
in this case we put $l=1$.
\end{sdefinition}

\begin{stheorem}{\bf Corollary.}
Let $X$ be a projective surface with \T-singularities and 
let  $\mathcal X/(0 \in S)$ be a $\QQ$-Gorenstein deformation 
induced by a local deformation of one point $P\in X$. 
Then, in the notation of \xref{Tdeformation}, for a general fiber $\mathcal X_s$, $s\in S$ we have
\[
\rho(\mathcal X_s)-\rho(X)=l-1,
\]

\end{stheorem}

Now we can take  one of the toric surfaces with \T-singularities
and $\rho(X)=1$ 
described in \cite[\S 4]{Hacking-Prokhorov-2010} 
and deform it in a suitable way.
\begin{mdefinition}{\bf Example.}
 Take the weighted projective plane $X:=\PP(a^2,b^2,5c^2)$,
 where $a$, $b$, $c$,  
are given by the following Markov-type equation 
\[
a^2+b^2+5c^2=5abc
\]
(cf. \cite{Karpov1998}).
Then $X$ has three singular points which are of type \T{} and $K_X^2=5$.
More precisely, 
\[
\Sing(X)=\left\{\frac1{a^2} (b^2,5c^2),\ \frac1{b^2}(a^2,5c^2), \  \frac1{5c^2}(a^2,b^2) \right\} 
\] 
For the third point we have 
\[
 \frac1{5c^2}(a^2,b^2)=\frac1{5c^2}(1,5c\alpha-1),
\]
where $\alpha=ab\delta$ and $\delta$ is taken so that $a^2\delta\equiv 1\mod 5c^2$.
Thus deforming this point  to one of the following collection of singularities
\begin{itemize}
\item  
$\frac1{c^2}(1,c\alpha-1)$,\ $A_3$;
\item  
$\frac1{2c^2}(1,2c\alpha-1)$,\ $A_2$;
\item  
$\frac1{3c^2}(1,3c\alpha-1)$,\ $A_1$;
\item  
$\frac1{4c^2}(1,4c\alpha-1)$,
\end{itemize}
we get examples  of 
del Pezzo surfaces as in (iii) 
of Theorem \ref{main} with $K_X^2=5$, $\rho(X)=2$, $\s(X)=3$.
\end{mdefinition}

\def\cprime{$'$} \def\mathbb#1{\mathbf#1} \def\bblapr{April}

\end{document}